A meeeting was held to discuss approaches to improving the performance of the

drive laser on the ITS. Attending were Shinn, Sinclair, Kehne, Bohn and Neil.

Identified problems include 1) Long term drift, 2) short term angular jitter, 3) possible movement of mounts caused by accidental personnel interference. To deal with #3 and probably also improve 1&2 several of the actions noted below are aimed at minimizing entry into the clean room.

A set of long term and short term fixes were identified, long term meaning during the shutdown or longer and short term means this week, perhaps today for many. The fixes were (by priority/schedule)

Short term:

1) Set up diffusers/curtains over the table to eliminate wind. Add additional shielding as necessary. This should minimize (but not eliminate) fast jitter

2) Disconnect z drive of telescope. There should never again be a change in relative clean room/cave spot size ratio. Check it occasionally (see #4) but leave it alone.

3) Send auto correlator signal to control room so pulse length can be monitored without entry into clean room.

4) Eliminate using Spirocon to view input beam by using ebeam itself in emittance dominated mode and solenoid to image electron emitting area onto harp to verify proper size. Model should give you exact measure of what spot is at cathode.

5) Replace remaining mounts in table transport with good ones. (Priority set by arrival date)

6) Remote the attenuator control to eliminate need for clean room access

7) Set up camera to monitor aperture image in clean room. This will verify proper input beam.

Long term:

1) Modify Brewster window to add input beam monitor.

2) Remote cavity length control

3) Remote bias and extinction ration optimization

4) (Remote adjust of laser table mirrors)

5) (Install Coherent Automatic Alignment System)

(4&5 to be done only if necessary based on results of measurements of performance of earlier fixes)

It was agreed that replacement of the apertures and spot size adjustment at the aperture would be done a seldom as possible since these must of necessity remain manual. It was agreed that we would determine the possibility of strictly top hat distributions with an overfilled aperture based on the QE and available power. This would eliminate most jitter but is questionable at the highest charges. Need better QE numbers from test. Based on the next two days experience we will evacuate the transport line to eliminate a remaining possible jitter source although no evidence to date suggests this is a problem.

All felt the above actions will go a long way toward not only giving good thesis data but put us closer to a usable FEL system.